OpenStudy (anonymous):

Can you tell me how to reengage this equation to get the right answer? for boiling point?

OpenStudy (anonymous):

\[\ln (\frac{ 1.50}{ 1 }) =-(\frac{ 40.7 }{ 8.03145 })(\frac{ 1 }{ x }-\frac{ 1 }{ 373.15 })\]